please help I am so confuse how to do this type of problem
Floor plans for a building have the four corners of a room located at then points (2, 3), (11, 6), (-3, 18), and (8, 21).determine whether the side
through the points (2, 3) and (11, 6) is perpendicular to the side through the points(2, 3) and (-3, 18).

Answer by stanbon(48558)

determine whether the side through the points (2, 3) and (11, 6) is perpendicular to the side through the points(2, 3) and (-3, 18).
----------
slope of the side thru (2,3), (11,6) is [6-3]/[11-2] = 3/9 = 1/3
slope of the side thru (2,3), (-3,18) is [18-3]/[-3-2] = 15/-5 = -3
-------------
Look at the product of the slopes (1/3)*-3 = -1
Since the product is -1 the sides are perpendicular.
===============
Cheers,
Stan H.